# GATE OpenING: A Comprehensive Benchmark for Judging Open-ended Interleaved Image-Text Generation

### v1

@@ 4.3 Training of IntJudge
To enhance the training of IntJudge, we propose a Reference-Augmented Generation (RAG) approach to scale up the training dataset. As illustrated in Fig. 3, we train our IntJudge model on the combination of human-annotated pairwise data from Dev Set and the RAG pairwise data. To generate scalable pairwise data using our RAG approach, models were provided with gold real-world answers from the Dev Set and prompted to generate responses based on these gold answers. A pairwise data is formed by pairing a plain generation result with an RAG-based output, with the RAG result assigned as the winner. A bag of models, including $g$ seen interleaved generation methods are used for plain generation and RAG. The total training objective is defined as:
@@ 5.1 Experimental Setup
Models. We test 10 representative methods and categorized them into three types: 1) Integrated pipeline consists of independent models for generating text and image in two stages, including GPT-4o+DALL-E·3 (DALL-E3) [48, 8] and Gemini1.5+Flux [63, 9]; 2) Two-stage generator uses a unified model architecture and outputs text and image in separate stages, including Emu2 [59], SEED-X [23], and Show-o [77]; 3) End-to-end generator outputs image-text results in a single stage, including GILL [35], NExT-GPT [74], MiniGPT-5 [88], SEED-LLaMA [22], and Anole [16]. We keep GPT-4o+DALL-E3, Anole, SEED-LLaMA, and NExT-GPT as unseen models for IntJudge validation, and the rest are models seen in IntJudge training.
@@
Evaluation Metrics. We assess model performance using two key metrics: win rate and agreement. Win rate indicates how often a model wins in pairwise comparisons. We consider four methods for handling ties: 1) Force Dividing Tie (FDT): We force judges to assign ties with a more leaning model in rules and prompts, ensuring that every comparison results in a decisive outcome. If a tie leans towards model A (Tie(A)), it counts as a win for A; similarly for B. This method allows for clear rankings without ambiguity. 2) Without Tie (w/o Tie): Tied comparisons are excluded; only matches with a clear winner are considered; 3) With Tie counted as 0 (w/ Tie (0)): Ties are included but do not contribute to the win count of either model; 4) With Tie counted as 0.5 (w/ Tie (.5)): Each tie contributes half a win to both models. Agreement measures the consistency between different evaluators (e.g., automated pipelines and human judgments) under the same tie-handling strategies. It reflects how often the evaluators concur in their assessments.
@@ 5.2 Overall Evaluation
Evaluation of Three Judges. We conduct experiments to evaluate the performance of different models using the win rate and agreement metrics. Table 2 showcases the win rates of various models under different judge methods, including Human, GPT-based, and IntJudge-based Evaluations. The sampling round $E$ is set in 2 to form 4,320 battle pairs. It is found that the integrated pipelines like GPT-4o+DALL-E3 and Gemini 1.5+Flux consistently perform better across all evaluators, while the end-to-end models like MiniGPT-5, GILL, and NExT-GPT exhibited lower performance.
@@
Pairwise Model Performance. We visualize the pairwise comparison results of all methods in Fig. 5, evaluated by human, GPT-4o, and IntJudge, respectively. The heat map reveals win-loss relationships, where warmer colors represent higher win rates and cooler colors vice versa. Notably, GPT-4o+DALL-E3 and Gemini1.5+Flux demonstrate strong win rates against other models, even comparable to the manually annotated output under GPT evaluation.
@@
Text-only and Image-only Evaluation. To explore whether text or image has a greater impact on the model performance, we evaluate models on text-only and image-only outputs on the same sampled pairs. Fig. 4 shows that MiniGPT-5 and GILL do not perform well mainly due to the low quality of text outputs. SEED-X and NExT-GPT achieve higher win rates on text-only evaluation, however, the lower quality of generated images hinders their opportunity to achieve higher ranking in Table 2. It is observed that text generated by GPT-4o even outperforms real-world content annotated by human, highlighting the superior language capabilities of GPT-4o.
GPT-based Scoring. As shown in Fig. 6, detailed scores from GPT-based evaluations are provided to support explainable performance analysis of different models. It is observed that GPT-4o+DALL-E3 performs suboptimally in meta-topics like Interactive Image Editing (IIE) and Embodied-AI Tasks (ET), possibly due to the lack of relevant training data in these scenarios. GPT-4o also showcases the inherent biases to its own generation results. For example, GPT-4o gives 10 scores to its own answers in Human Preference Alignment. In contrast, outputs annotated by humans only achieve an average score of 9 in human preference alignment.
@@ 5.3 Ablation Studies
Ablation on Judge Training Data. We investigate the influence of incorporating RAG data on the performance of the IntJudge. The comparison is conducted between two training configurations: one utilizing only the arena data (6,014 samples) and the other augmented with RAG data (25,982 samples). As illustrated in Fig. 8, with RAG data included, the FDT agreement on unseen models increases by 7.8%, demonstrating the effectiveness of our RAG-based strategy.
@@
Ablation on Image Generator. We sample 200 data instances across tasks to assess the influence of different image generators on interleaved performance. Table 4 presents a comparison of several basic text generation methods combined with different image generators. We assign the text generated by Gemini with Flux-schnell (Flux-sch) and Flux-dev to explore the difference. The results highlight that advanced image generators like Flux-dev are crucial for improving the overall quality of interleaved content. Meanwhile, it is noted that the generation efficiency of Flux-dev is significantly slow. This prompts further exploration of flow models to better balance generation quality and efficiency.
@@
Figure 9: Error distribution of three models: GPT-4o+DALL-E3 (integrated), SEED-X (two-stage), and Anole (end-to-end).
@@ 5.4 Analysis and Discussions
Error Analysis. We conduct an error analysis on 200 sampled instances where the output of models performed worse than human output. Fig. 9 shows the frequency of error types across three different types of models, providing insights into their specific performance limitations. GPT-4o+DALL-E3 suffers from content incoherency and inconsistency since it is hard for DALL-E3 to generate multiple images in the same style. Poor image quality is the main problem of Anole, as its finetuning data for image generation is insufficient. While most outputs by SEED-X have multiple errors, the inexistence of text or image content is still the major problem.
No-Image and No-Text Ratios. Table 5 presents the no-image, no-text, and no-image-and-text ratios of different models, representing the proportion of instances where the models fail to generate visual content, textual content, and both. Human, GPT-4o+DALL-E3, and Genimi1.5+Flux exhibited near-zero failure rates (except that image generation on certain sensitive topics is prohibited according to content safety policies), indicating consistent multimodal generation. On the other hand, models like SEED-X and NExT-GPT showed a high no-image ratio. This can be attributed to poor instruction following and generation ability. These findings suggest that models need to generate both images and text properly to achieve higher rankings on OpenING.
@@
Findings and Discussions. We discuss key findings from our experiments to inspire future works: 1) While all generative models ranked lower than Human in interleaved generation, end-to-end models still lagged significantly behind integrated pipelines like GPT-4o+DALL-E3 in interleaved generation tasks. Two-stage generation methods based on unified models are also expected to be further improved. 2) Natural images consistently outperform generated images, highlighting the challenge of generating high-quality images. 3) Text generated by GPT can match or even surpass the quality of human-annotated text, showcasing the effectiveness of LLMs in producing rich and informative content. 4) Foundational generative models play a big role. For example, when combining GPT, Gemini, and SEED-X with Flux-dev, respectively, all models demonstrate higher performance. 5) As large-scale data is crucial for training judge models, using our RAG method to scale up data beyond manual annotation contributed to the building of a more robust judge model.
@@ 6 Conclusion
We introduce OpenING, a comprehensive benchmark designed to evaluate open-ended interleaved image-text generation. OpenING overcomes the limitations of existing benchmarks by including more diverse data and tasks grounded in real-world scenarios. To address the challenges in assessing open-ended multimodal generation, we propose a robust judge model IntJudge, which is trained on both human-annotated and RAG-based data from the Dev Set of OpenING. We evaluate various interleaved generation methods on the Test Set of OpenING, revealing the significant challenges for generating coherent and high-quality interleaved image-text content. Ablation studies further demonstrate the effectiveness of our RAG-based data pipeline in training IntJudge. Looking forward, increasing the size and diversity of the data could further enhance the practical relevance of interleaved generation benchmarks. We anticipate that OpenING will inspire further research in MLLMs, and support the development of multimodal evaluation models.
@@ A.1 Hierarchical Structure of OpenING
We present in Table 6 all 56 specific tasks from OpenING, belonging to 23 meta-topics, including the number of Instances (# of Ins.), Meta-topic names, and capabilities of MLLMs evaluated in testing. Specifically, the 56 tasks are divided into 38 common tasks and 18 hard tasks. 38 common tasks are annotated by 28 professional annotators, instructed and monitored by 14 data experts. 18 hard tasks are annotated by 14 data experts as these tasks need certain domain knowledge for data extraction, collection, and annotation. It is noted that 18 hard tasks are extremely difficult, especially in collecting data from rare sources and annotating data into the desired format. The number of instances in each category is also provided.
@@ A.2 Data Sources
We list all sources where meta-data are collected for the annotation of our OpenING benchmark. Annotators arrange the images collected from the sources into the standardized multi-step format and annotate the corresponding text for each image. The details of the data source are presented in Table LABEL:tab:data_source, including the ID number of each task. We also provide examples of the source of data instances to show how the desired data are searched from a certain platform.
A.3 Task Prompt Breakdown
We present a comprehensive breakdown of the task prompts used in our experiments. The tasks are designed to evaluate various capabilities of interleaved image-text generation methods, ranging from storytelling and creative design to problem-solving and interactive experiences. For each task, we provide the general prompt format and specific prompt examples to illustrate how the models are expected to generate interleaved image-text outputs in Table LABEL:tab:prompt_examples. By carefully designing the general prompt templates for these tasks and refining their corresponding prompt examples in data instances we aim to challenge all interleaved generation methods using more diverse queries, showing the generalization performance of methods more comprehensively.
A.4 Exclusive Protocols for Data Filtering
@@
A.5 Task Abbreviations
Given the large number of tasks and methods tested in our benchmark, we use abbreviations for simplification in the manuscript. In order to facilitate understanding of possible abbreviations in the text, we provide respectively the abbreviations of the 23 meta-topics in Table 8 and the abbreviations of the 56 tasks in Table 9.
A.6 Interface of IntLabel
To ensure that annotations were consistent within the standard format across annotators, and to facilitate the annotation process, we developed a labeling tool for interleaved image-text data annotation, named IntLabel. All queries and answers in OpenING are annotated manually, checked and refined via our developed tool, IntLabel. The final saving format of the annotation is JSONL, alongside a folder containing the corresponding image. We present the GUI of our IntLabel in Fig. 10. IntLabel is developed based on PyQt [^5] https://www.riverbankcomputing.com/software/pyqt. We will also release the source code of IntLabel for contributing to the community.
A.7 Illustration of More Data Examples
@@
Illustrations of the representative example of the 23 meta-topics are provided in Fig. 18 to showcase the diversity and complexity of tasks in our benchmark. These examples highlight the variety of interleaved image-text data that OpenING encompasses, demonstrating the challenges and capabilities required for effective interleaved image-text generation.
@@
Table 7: Task Sources: ID number and data sources of each task.
@@ 18.2 Chat with Memes MOD [21] and in-house conversations
23.1 Autonomous Driving and In-door Navigation CARLA [19], VisDrone [93], Gibson Environment (http://gibsonenv.stanford.edu/) and Reverie (https://reverie.herokuapp.com/arXiv_Demo/)
@@
Abbrev. Meta-Topic Name
SC Storybook Creation
MRG Multimodal Report Generation
MCC Multimodal Content Completion
MLG Multimodal Layout Generation
GN GUI Navigation
IIE Interactive Image Editing
IVD Interactive Visual Design
ME Multimodal Exam
GG Graph Generation
ER&DS Event Reasoning & Deductive Simulation
2IR 2D Image Reasoning
@@
MC Multimodal Information Summary
IR Multimodal Information Recommendation
MB Multimodal Brainstorming
TSF Multimodal Time Series Forecasting
GT Geographical Tasks
SMT Social Media Tasks
FT Fashion Tasks
CT Cooking Tasks
ET Educational Tasks
HT Healthcare Tasks
EAT Embodied-AI Tasks
Table 9: Abbreviations of Tasks. Each task abbreviation is followed by its full term for clear reference.
Abbrev. Task Name Abbreviation Task Name

### v3

@@ 4.3 Training of IntJudge
To enhance IntJudge training, a Reference-Augmented Generation (RAG) approach is proposed to scale up the training data set. As shown in Fig. 3(b), our IntJudge model is trained on the combination of human-annotated pairwise data from the Dev Set and the RAG pairs. In our RAG approach, models are provided with gold real-world answers from the Dev Set and prompted to generate responses based on these gold answers. Pairwise data are formed by pairing a plain generation result with an RAG-based output, where the RAG result is assigned as the winner. A bag of models, including $g$ seen interleaved generation methods are used for plain generation and RAG. The training objective is defined as:
@@ 5.1 Experimental Setup
[…]
@@
Evaluation Metrics. Model performance are evaluated using two key metrics: win rate and agreement. Win rate indicates how often a model wins in pairwise comparisons. Four methods used to handle ties include 1) Force Dividing Tie (FDT): We force judges to assign ties with a more leaning model in rules and prompts, ensuring that every comparison round results in a decisive outcome. A win is attributed to A if a tie favors model A (Tie(A)), likewise for B. This metric allows for clear rankings without ambiguity. 2) Without Tie (w/o Tie): Tied comparisons are excluded; only matches with a clear winner are considered; 3) With Tie counted as 0 (w/ Tie (0)): Ties are included but do not contribute to the win count of either model; 4) With Tie counted as 0.5 (w/ Tie (.5)): Each tie contributes half a win to both models. Agreement measures the consistency between different evaluators (e.g., automated pipelines and human judgments) under the same tie-handling strategies. It tells the frequency with which the evaluators agree in their assessments.
@@ 5.2 Overall Evaluation
[…]
@@
[…]
@@
Text-only and Image-only Evaluation. To explore the impact of text and image on model performance, we evaluate models using text-only and image-only outputs on the same sampled pairs. Fig. 4 shows that MiniGPT-5 and GILL underperform primarily due to the low quality of their text outputs. SEED-X and NExT-GPT achieve higher win rates on text-only evaluation, however, the lower quality of generated images limits their ranking as shown in Table 2. Text generated by GPT-4o even outperforms human-annotated content, which demonstrates its superior language capabilities.
[…]
@@ 5.3 Ablation Studies
Ablation on Judge Training Data. We investigate the influence of incorporating RAG data on the performance of the IntJudge. The comparison is conducted between two training configurations: one that utilizes only the arena data (6,014 samples), the other being augmented with RAG data (25,982 samples). As illustrated in Fig. 8, with RAG data included, the FDT agreement on unseen models increases by 7.8%, demonstrating the effectiveness of our RAG-based strategy.
@@
Ablation on Image Generator. We sample 200 data instances from all tasks to assess the influence of image generators on interleaved performance. Table 4 compares basic text generation methods paired with different image generators. The results suggest that image generators greatly affect the quality of interleaved generation. For example, performance improves greatly when text models are paired with Flux-dev over other image models. It is also noted that Flux-dev has slower generation efficiency despite better image quality than Flux-schnell.
@@
[…]
@@ 5.4 Analysis and Discussions
[…]
@@
Findings and Discussions. We discuss key findings from our experiments to inspire future works: 1) All generative models ranked lower than human in interleaved generation. The unified end-to-end models lagged significantly behind integrated pipelines, which combine more developed foundation models. The unified two-stage generation methods also need further improvements. 2) Natural images consistently outperform generated ones, indicating the significant challenges of high-quality image generation. 3) The quality of GPT-generated text can be comparable to or even exceed that of human-annotated text, demonstrating the effectiveness of LLMs in producing rich and informative textual content. 4) Image generation has a great impact on interleaved generation. The quality of interleaved content improves markedly when text models are paired with more advanced image models. 5) Large-scale data is crucial for training judge models. By scaling up data beyond manual annotations, our RAG method contributed to training a more robust judge model.
@@ 6 Conclusion
We introduce OpenING, a comprehensive benchmark for evaluating open-ended interleaved image-text generation. OpenING addresses the limitations of existing benchmarks by a wider coverage of more diverse data and tasks grounded in real-world scenarios. To better assess open-ended multimodal generation, we propose IntJudge, a robust judge model trained on both human-annotated and RAG-based data from the Dev Set of OpenING. It is expected that our IntJudge can serve as a reward model in future RL-based (e.g., GRPO) generative models. Evaluation of various interleaved generation methods on the Test Set of OpenING reveals the challenges of generating coherent and high-quality interleaved image-text content. Ablation studies reaffirm the effectiveness of our RAG-based data pipeline for training IntJudge. Looking ahead, expanding the size and diversity of interleaved generation benchmarks could unlock even greater real-world potential and impact. We anticipate that our OpenING will inspire future research in MLLMs and benefit the development of multimodal evaluation models.
OpenING: A Comprehensive Benchmark for Judging
Open-ended Interleaved Image-Text Generation
Supplementary Material
In this supplementary material, we provide additional information, discussions, and results in support of the primary text, which are organized as follows:
Sec. A includes the details of OpenING and data curation.
Sec. B presents the details of IntJudge.
Sec. C presents the details of evaluation.
Sec. D provides the details of our experiments and additional evaluation results.
Sec. E introduces the extended experiments and results on fine-tuning MLLMs using our OpenING benchmark.
Sec. F discusses the limitations of this study.
The indexings of all figures, tables, and equations continue from the primary text.
@@ A.1 Hierarchical Structure of OpenING
We present in Table 6 the 56 specific tasks from OpenING that derived from 23 meta-topics. The number of Instances (# of Ins.), Meta-topic names, and capabilities of MLLMs evaluated in testing are provided. Based on the required annotation skills, the 56 tasks are divided into 38 common tasks and 18 hard tasks. The common tasks are annotated by 28 professional annotators, instructed and supervised by 14 data experts. The annotations of hard tasks, which require specific domain knowledge and special data reasoning and processing techniques, are conducted by the 14 data experts.
@@ A.2 Data Sources
We list all sources where meta-data are collected for the annotation of our OpenING benchmark. Annotators arrange the images collected from the sources into the standardized multi-step format and annotate the text for each corresponding image. The details of the data source are presented in Table LABEL:tab:data_source, including the ID number of each task. We also provide examples of the source of data instances to show how the desired data are searched from a certain platform.
A.3 Exclusive Protocols for Data Filtering
@@
A.4 OpenING Data
@@
Illustrations of the representative example of the 23 meta-topics are provided in Fig. 18 to showcase the diversity and complexity of tasks in our OpenING benchmark. These examples highlight the variety of interleaved image-text data that OpenING encompasses, demonstrating the challenges and capabilities required for effective interleaved image-text generation.
A.5 Annotation of OpenING
To facilitate the annotation process and ensure consistent annotations standards across annotators, we developed and opensourced an annotation tool for interleaved image-text data labeling. Our IntLabel annotation tool is developed based on PyQt [^5] https://www.riverbankcomputing.com/software/pyqt, and the IntLabel GUI is presented in Fig. 10. All queries and answers in OpenING are annotated, checked, and refined manually via IntLabel. The annotated JSONL files and corresponding images are kept in the same folder.
A.6 Task Abbreviations
Given the large number of tasks in our benchmark and the wide range of methods evaluated, the abbreviations of the 23 meta-topics and 56 tasks are provided respectively in Table 8 and Table 9.
Figure 11: Annotation Interface of Interleaved Arena for human judges to compare the anonymous outputs of model A and model B. Human evaluators are instructed to select "A is Better" or "B is Better" to choose a winner for the pairwise comparison. When the two outputs are similar in quality and a tie option has to be chosen, human evaluators are instructed to select "Tie (A is slightly Better)" (Tie(A)) or "Tie (B is slightly Better)" (Tie(B)). Zoom in for a better experience.
Appendix B Details of IntJudge and Arena Data
B.1 Annotation of Interleaved Arena
The Interleaved Arena is introduced as an evaluation framework specifically to address the challenges of assessing open-ended interleaved image-text generation. Evaluating interleaved image-text generation is difficult due to: 1) Multiple images and text may need to be assessed at the same time; 2) The answer may be open-ended, where there is no single correct answer (multiple solutions exist for a query). Comparative evaluation has been demonstrated to offer greater stability and reliability than subjective scoring [14, 93, 38]. Therefore, pairwise comparison is employed to ensure consistency and accuracy. The Interleaved Arena facilitates this by supporting evaluations from human judges, GPT-based judges, and the proposed IntJudge. The Interleaved Arena consists of two main components: 1) A sampling strategy to select pairwise comparison from all available interleaved generation methods fairly, and 2) An annotation interface for human judges to conduct evaluations manually. As shown in Fig. 11, the annotation interface is developed using PyQt and will be made available to researchers for facilitating future development. Using the annotation interface of Interleaved Arena, annotators are tasked with comparing anonymous outputs from two multimodal agents for each input prompt and deciding which is the winner based on seven predefined criteria. The vote results are used to rank interleaved generation models based on their win rates in Interleaved Arena. Since the previous studies [93, 14] noted that too many ties cause inefficiency, our annotators are instructed to appoint a more leaning output when choosing a tie for a battle pair, denoted as Tie(A) or Tie(B).
B.2 Training of IntJudge
The IntJudge’s architecture is based on Qwen2-VL-7B. Hyperparameters include cutoff length 16,240, batchsize 8, learning rate 1.0e-4, training epochs 20, training precision BF16, and training dataset (31,996 samples, with 6,014 arena data and 25,982 RAG data). As defined in the Eq. 7 in the primary text (see Section 4.7), the total loss consists of cross-entropy loss $\mathcal{L}_{\text{CE}}$, contrastive loss $\mathcal{L}_{\text{CT}}$, MSE loss $\mathcal{L}_{\text{MSE}}$, and pairwise ranking loss $\mathcal{L}_{\text{PR}}$:
where the loss weights are $\lambda_{1}=1.0$, $\lambda_{2}=0.01$, $\lambda_{3}=0.01$, $\lambda_{4}=0.01$ respectively.
Cross-entropy loss for language modeling is given by:
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{N}y_{i}\log(\hat{y}_{i}),$ (8)
Contrastive Loss aligns image and text embeddings, which is written as:
Mean Squared Error Loss for image feature regression and reducing prediction errors is given by:
Pairwise Ranking Loss for precise rankings of outputs weights the relative quality of generated outputs in pairwise comparison, which is given by:
Appendix C Details of Evaluation
C.1 Prompt for Interleaved Generation
Each interleaved generation prompt is carefully constructed by combining two main parts: task-specific prompts and model-specific prompts. Model-specific prompts are adjusted based on the individual characteristics of each model, providing hints and guidance to ensure coherent and consistent outputs across multiple steps, as illustrated in Table LABEL:tab:model_prompt. In our proposed Reference Augmented Generation (RAG) pipeline, certain seen models are prompted with gold answers to produce more precise outputs.
C.2 Task Prompt Breakdown
We present a comprehensive breakdown of the task prompts used in our experiments. The tasks chosen are diverse, aiming to thoroughly test various capabilities of interleaved image-text generation models. Specifically, these tasks range from storytelling and creative design to problem-solving and interactive experiences. For every task, we outline the general format of the prompt and present specific examples, as shown in Table LABEL:tab:prompt_examples. These detailed examples demonstrate clearly how models are expected to produce combined image and text sequences. By carefully designing the general prompt templates for these tasks and refining their corresponding prompt examples in data instances, we aim to challenge all interleaved generation methods using more diverse queries. This comprehensive approach allows us to assess the generalization abilities of different interleaved generation methods more effectively.
@@ C.2 Task Prompt Breakdown
Table 7: Task details: ID number, name, and data sources of each task.
@@ 18.2 Chat with Memes MOD [21] and in-house conversations
23.1 Autonomous Driving and In-door Navigation CARLA [19], VisDrone [97], Gibson Environment (http://gibsonenv.stanford.edu/), Reverie (https://reverie.herokuapp.com/arXiv_Demo/), and Matterport 3D (https://aihabitat.org/datasets/hm3d/)
@@
Abbrev. Meta-Topic Name Abbrev. Meta-Topic Name
SC Storybook Creation MC Multimodal Information Summary
MRG Multimodal Report Generation IR Multimodal Information Recommendation
MCC Multimodal Content Completion MB Multimodal Brainstorming
MLG Multimodal Layout Generation TSF Multimodal Time Series Forecasting
GN GUI Navigation GT Geographical Tasks
IIE Interactive Image Editing SMT Social Media Tasks
IVD Interactive Visual Design FT Fashion Tasks
ME Multimodal Exam CT Cooking Tasks
GG Graph Generation ET Educational Tasks
ER&DS Event Reasoning & Deductive Simulation HT Healthcare Tasks
2IR 2D Image Reasoning EAT Embodied-AI Tasks
@@
Table 9: Abbreviations of Tasks. Each task abbreviation is followed by its full term.
Abbrev. Task Name Abbrev. Task Name
@@
C.3 Key Evaluation Criteria
  1. 1)
  2. 2)
  3. 3)
  4. 4)
  5. 5)
    Human Preference Alignment: Outputs are evaluated to ensure they align with general human preferences, avoiding offensive, inappropriate, or misleading contents.
  6. 6)
  7. 7)
C.4 Prompt for GPT-based Judge
The prompt used for a GPT-based judge is illustrated in Fig. 12, where GPT-4o is employed to compare the quality of answers generated by two interleaved generation methods (Model A vs. Model B) for a given input question. The evaluation is based on seven criteria: Correctness, Image-Text Coherency, Multi-step Consistency, Content Quality, Human Preference Alignment, Completeness, and Content Richness. The judge is required to compare the overall quality of the responses, determine which model performed better, and output a clear verdict (e.g., "A is better," "B is better"). The judge must choose a more favorable model if a Tie case is determined. This hierarchical judging approach allows for a thorough criterion-driven comparison of the two generated answers, contributing to a detailed understanding of the relative strengths of each model. Judges models based on Qwen2-VL and InternLM-XComposer2.5 were explored using the same system prompt. In order to reduce the number of input tokens when implementing these open-source MLLMs, we further refined these system prompts.
C.5 Prompt for Qwen, Intern and IntJudge
System prompts are designed for optimal judgments based on Qwen2-VL, InternLM-XComposer2.5 and our IntJudge (see Fig. 13). These prompts were refined through extensive prompt engineering to maximize efficiency and reduce token usage, and ultimately reduce GPU memory usage in evaluating the open-source MLLMs. The prompt instructs the models to compare the quality of answers generated by two methods, named Model A and Model B. The goal of the design is to provide an objective assessment that aligns with human evaluators. We also provide a previously unrefined prompt for Qwen2-VL and InternLM-XComposer2.5 for comparison in Fig. 13. The refined prompt allows for more streamlined input, ensuring the judgments are concise while still covering all essential evaluation aspects.
C.6 Prompt for GPT-based Scoring
The system prompt designed for GPT-based evaluators is presented in Fig. 14. The prompt instructs GPT to evaluate interleaved image-text content based on seven key criteria. For each criterion, the GPT-based evaluator rates the model on a scale from 0 to 10 and provides a brief explanation to justify the assessment. The GPT-based evaluations serve as a supplementary analysis of the generated content, supporting further performance comparisons between models.
Appendix D Details of Experiments
D.1 Baseline Methods
In this section, we provide more details of the 12 representative methods we evaluated in the primary text. These methods are categorized into three groups: 1) Integrated pipeline, which involves separate models for text and image generation in two stages, such as GPT-4o+DALL-E·3 (DALL-E3) [48, 8] and Gemini1.5+Flux[64, 9]; 2) Two-stage generator, which employs a unified model architecture to produce text and images in separate stages, including Emu3 [71], VILA-U [77], Emu2 [60], SEED-X [23], and Show-o [79]; and 3) End-to-end generator, which directly generates image-text outputs in a single step, such as GILL [35], NExT-GPT [75], MiniGPT-5 [92], SEED-LLaMA [22], and Anole [16]. For IntJudge validation, we reserve GPT-4o+DALL-E3, Emu3, VILA-U, Anole, SEED-LLaMA, and NExT-GPT as unseen models for IntJudge validation, while the remaining models are regarded as seen models and included in IntJudge training.
The advantage of the Integrated Generation Pipeline lies in its modularity, allowing each component to specialize in its respective task—text generation or image creation. This approach leverages the strengths of SOTA proprietary models like GPT-4o and DALL·E-3 to produce coherent and visually compelling interleaved outputs. However, its two-stage nature may introduce latency and potential alignment challenges between text and images. Similarly, combining Gemini 1.5 with Flux benefits from the robust text generation capabilities of Gemini 1.5 and the efficient image generation of Flux-schnell. This setup enables high-quality content production while maintaining the flexibility of modular design. Nevertheless, synchronization and contextual consistency between pipeline stages still need improvement.
Two-stage Interleaved Generator, Emu2, SEED-X and Show-o are implemented to output text and image in two stages based on a unified model architecture. We also introduce two of the latest models: Emu3 and VILA-U. Emu3 improves Emu2 by training entirely on next-token prediction, capable of generating more high-quality images, videos, and text by tokenizing multimodal sequences into a discrete space and training a single transformer. Similarly, VILA-U can work as a two-stage approach through a single autoregressive next-token prediction framework, enabling precise alignment and increased fidelity in multimodal content.
[…]
Model Prompt Format
Integrated Pipelines
[…]
Two-Stage Generators
Emu2 Instruction = $\big{[}\langle T_{i}\rangle+\langle\mathrm{img_{in,i}}\rangle$ for $i=1,\cdots,M\big{]}$. Loop $j=1,\cdots,N$: Text_Answersj = Emu2.TextGen(Instruction); Image_Answersj = Emu2.ImgGen(“The prompt for this generation is:" + Text_Answersj + “The context of this task is:" + Instruction)
[…]
Emu3 Instruction = $\big{[}\langle T_{i}\rangle+\langle\mathrm{img_{in,i}}\rangle$ for $i=1,\cdots,M\big{]}$. Loop $j=1,\cdots,N$: Text_Answersj = Emu3.TextGen(Instruction); Image_Answersj = Emu3.ImgGen(“The prompt for this generation is:" + Text_Answersj + “The context of this task is:" + Instruction)
SEED-X Instruction = $\big{[}\langle\mathrm{img_{in,i}}\rangle$ for $i=1,\cdots,M\big{]}$ + $\big{[}\langle T_{i}\rangle$ for $i=1,\cdots,M\big{]}$. Loop $j=1,\cdots,N$: Text_Answersj = SEED-X.TextGen(Instruction); Image_Answersj = SEED-X.ImgGen(Instruction + Text_Answersj)
[…]
Show-o Instruction = [$\big{[}\langle T_{i}\rangle+\langle\mathrm{img_{in,i}}\rangle$ for $i=1,\cdots,M\big{]}$. Loop $j=1,\cdots,N$: Text_Answersj = Show-o.TextGen(Instruction); Image_Answersj = Show-o.ImgGen(“The prompt for this generation is:" + Text_Answersj + “The context of this task is:" + Instruction)
[…]
Vila-U Instruction = $\big{[}\langle T_{i}\rangle+\langle\mathrm{img_{in,i}}\rangle$ for $i=1,\cdots,M\big{]}$]. Loop $j=1,\cdots,N$: Text_Answersj = Vila-U.TextGen(Instruction); Image_Answersj = Vila-U.ImgGen(Text_Answersj + Instruction)
End-to-End Generators
[…]